OpenStudy (anonymous):

AWARDING MEDALS! F(x) = (x-1) (x-2) includes 2 real rational solutions because?

OpenStudy (anonymous):

x=2 and x=1 are real an rational numbers?

OpenStudy (anonymous):

If you are asking me if that is the answer, I don't know. I was asking because I have to know

OpenStudy (anonymous):

irrational numbers are \[\pi, \sqrt{2}, etc.. \] not whole numbers means the exact values are unending decimals. real numbers are whole numbers, fractions of 2 integers, with the denominator not equals to 0. so x=1,x=2 must be real and rational solutions. real is anything \[(-\infty,\infty)\]

OpenStudy (anonymous):

^ sorry i mean rational numbers, are whole numbers, fractions, cont..

OpenStudy (anonymous):

oh okay! That is much clearer to me now. I the different types of numbers are confusing to me sometimes